A certain reaction is second order in N2 and first order in H2. Use this information to complete the table below. Round each of your answers to 3 significant digits.

[N2] [H2] initial rate of reaction
0.229 M 2.33 M 57.0 M/s
0.111 M 2.33 M __________M/s
0.0482 M 11.1 M ___________M/s
Answer #1

Rate of a reaction= K[A]*m[B]*n where K is rate constant,
[A],[B] are concentrations of reactants and the powers m and n are
order of reactants

here Rate = k[N2]*2[H2}*1 since N2 is 2nd order and H2 is first order

LETS see first row and find out rate constant k

57= k x (0.229)*2 x(2.33)*1

k= 57/0.122 =467.2M*(-2)S(-1)

this rate constant value can be used to fill next rows

2nd row
Rate= 467.2 x [0.111}*2x 2.33 =13.41 M/s

3rd row

Rate= 467.2x (0.0482)*2x 11.1=12.048 M/s

THANK YOU!